OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66.

OpenStudy (callisto):

\[a_n = a_1 + (n-1)d\]where n is the nth term, and d is the common difference. First find the common difference. \[a_{27} = a_1 + (27-1)d\]\[66= -12 + (26)d\]Solve d. Then, sub. n=42, \(a_1=-12\) and the value of d you've found into the formula to find the 42nd term. \[a_{42} = -12 + (42-1)d = ...\]

OpenStudy (anonymous):

Thanks! so, if d=3, a_42= 111? Thanks so much!! do you think you could help me answer another similar question?

OpenStudy (callisto):

Yup. :)

OpenStudy (anonymous):

Identify the 27th term of an arithmetic sequence where a1 = 38 and a17 = -74. uh and also Identify the 31st term of an arithmetic sequence where a 1 = 26 and a22 = -226. :) thanks again!

OpenStudy (callisto):

Please post a new question in a new post next time. Thanks! Btw, can you try to do it first? Use the above formula, plug in the suitable values and solve it.

OpenStudy (anonymous):

oh ok!!! sorry i am new lol, still getting used to the layout/rules

OpenStudy (callisto):

Welcome to OpenStudy! You're recommended to read the Code of Conduct at http://www.openstudy.com/code-of-conduct when you have time. Btw, what have you got for finding the 27th term question?

OpenStudy (anonymous):

-74=38+(17-1)d -112=16d d=-7 and then....

OpenStudy (callisto):

a27 = a1 + (n-1)d = ...?

OpenStudy (anonymous):

so a_27= -144?

OpenStudy (callisto):

Yup :) You can try the next one now, can you?

OpenStudy (anonymous):

yeah, i think i got it! thank you so much for all your help, i was not getting this concept at all!

OpenStudy (callisto):

Welcome :)